$SPAD/src/input poly1.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool poly1.output
)set message test on
)set message auto off
)clear all
--S 1 of 46
x + 1
--R
--R
--R   (1)  x + 1
--R                                                    Type: Polynomial(Integer)
--E 1
--S 2 of 46
z - 2.3
--R
--R
--R   (2)  z - 2.3
--R                                                      Type: Polynomial(Float)
--E 2
--S 3 of 46
y**2 - z + 3/4
--R
--R
--R               2   3
--R   (3)  - z + y  + -
--R                   4
--R                                          Type: Polynomial(Fraction(Integer))
--E 3
--S 4 of 46
y **2 + x*y + y
--R
--R
--R         2
--R   (4)  y  + (x + 1)y
--R                                                    Type: Polynomial(Integer)
--E 4

--S 5 of 46
% :: DMP([y,x],INT)
--R
--R
--R         2
--R   (5)  y  + y x + y
--R                       Type: DistributedMultivariatePolynomial([y,x],Integer)
--E 5
--S 6 of 46
p := (y-1)**2 * x * z
--R
--R
--R            2
--R   (6)  (x y  - 2x y + x)z
--R                                                    Type: Polynomial(Integer)
--E 6
--S 7 of 46
q := (y-1) * x * (z+5)
--R
--R
--R   (7)  (x y - x)z + 5x y - 5x
--R                                                    Type: Polynomial(Integer)
--E 7
--S 8 of 46
factor(q)
--R
--R
--R   (8)  x(y - 1)(z + 5)
--R                                          Type: Factored(Polynomial(Integer))
--E 8
--S 9 of 46
p - q**2
--R
--R
--R   (9)
--R         2 2     2     2  2          2      2       2             2
--R     (- x y  + 2x y - x )z  + ((- 10x  + x)y  + (20x  - 2x)y - 10x  + x)z
--R   +
--R          2 2      2       2
--R     - 25x y  + 50x y - 25x
--R                                                    Type: Polynomial(Integer)
--E 9
--S 10 of 46
gcd(p,q)
--R

--R
--R   (10)  x y - x
--R                                                    Type: Polynomial(Integer)
--E 10
--S 11 of 46
factor %
--R
--R
--R   (11)  x(y - 1)
--R                                          Type: Factored(Polynomial(Integer))
--E 11
--S 12 of 46
lcm(p,q)
--R
--R
--R             2             2        2
--R   (12)  (x y  - 2x y + x)z  + (5x y  - 10x y + 5x)z
--R                                                    Type: Polynomial(Integer)
--E 12
--S 13 of 46
content p
--R
--R
--R   (13)  1
--R                                                        Type: PositiveInteger
--E 13
--S 14 of 46
resultant(p,q,z)
--R
--R
--R           2 3      2 2      2      2
--R   (14)  5x y  - 15x y  + 15x y - 5x
--R                                                    Type: Polynomial(Integer)
--E 14
--S 15 of 46
resultant(p,q,x)
--R
--R
--R   (15)  0
--R                                                    Type: Polynomial(Integer)
--E 15
--S 16 of 46
mainVariable p
--R

--R
--R   (16)  z
--R                                                      Type: Union(Symbol,...)
--E 16
--S 17 of 46
mainVariable(1 :: POLY INT)
--R
--R
--R   (17)  "failed"
--R                                                    Type: Union("failed",...)
--E 17
--S 18 of 46
ground? p
--R
--R
--R   (18)  false
--R                                                                Type: Boolean
--E 18
--S 19 of 46
ground?(1 :: POLY INT)
--R
--R
--R   (19)  true
--R                                                                Type: Boolean
--E 19
--S 20 of 46
variables p
--R
--R
--R   (20)  [z,y,x]
--R                                                           Type: List(Symbol)
--E 20
--S 21 of 46
degree(p,x)
--R
--R
--R   (21)  1
--R                                                        Type: PositiveInteger
--E 21
--S 22 of 46
degree(p,y)
--R
--R
--R   (22)  2

--R                                                        Type: PositiveInteger
--E 22
--S 23 of 46
degree(p,z)
--R
--R
--R   (23)  1
--R                                                        Type: PositiveInteger
--E 23
--S 24 of 46
degree(p,[x,y,z])
--R
--R
--R   (24)  [1,2,1]
--R                                               Type: List(NonNegativeInteger)
--E 24
--S 25 of 46
minimumDegree(p,z)
--R
--R
--R   (25)  1
--R                                                        Type: PositiveInteger
--E 25
--S 26 of 46
totalDegree p
--R
--R
--R   (26)  4
--R                                                        Type: PositiveInteger
--E 26
--S 27 of 46
leadingMonomial p
--R
--R
--R            2
--R   (27)  x y z
--R                                                    Type: Polynomial(Integer)
--E 27
--S 28 of 46
reductum p
--R
--R
--R   (28)  (- 2x y + x)z
--R                                                    Type: Polynomial(Integer)

--E 28
--S 29 of 46
p - leadingMonomial p - reductum p
--R
--R
--R   (29)  0
--R                                                    Type: Polynomial(Integer)
--E 29
--S 30 of 46
leadingCoefficient p
--R
--R
--R   (30)  1
--R                                                        Type: PositiveInteger
--E 30
--S 31 of 46
p
--R
--R
--R             2
--R   (31)  (x y  - 2x y + x)z
--R                                                    Type: Polynomial(Integer)
--E 31
--S 32 of 46
eval(p,x,w)
--R
--R
--R             2
--R   (32)  (w y  - 2w y + w)z
--R                                                    Type: Polynomial(Integer)
--E 32
--S 33 of 46
eval(p,x,1)
--R
--R
--R           2
--R   (33)  (y  - 2y + 1)z
--R                                                    Type: Polynomial(Integer)
--E 33
--S 34 of 46
eval(p,x,y**2 - 1)
--R
--R
--R           4     3

--R   (34)  (y  - 2y  + 2y - 1)z
--R                                                    Type: Polynomial(Integer)
--E 34
--S 35 of 46
D(p,x)
--R
--R
--R           2
--R   (35)  (y  - 2y + 1)z
--R                                                    Type: Polynomial(Integer)
--E 35
--S 36 of 46
D(p,y)
--R
--R
--R   (36)  (2x y - 2x)z
--R                                                    Type: Polynomial(Integer)
--E 36
--S 37 of 46
D(p,z)
--R
--R
--R            2
--R   (37)  x y  - 2x y + x
--R                                                    Type: Polynomial(Integer)
--E 37
--S 38 of 46
integrate(p,y)
--R
--R
--R          1    3      2
--R   (38)  (- x y  - x y  + x y)z
--R          3
--R                                          Type: Polynomial(Fraction(Integer))
--E 38
--S 39 of 46
qr := monicDivide(p,x+1,x)
--R
--R
--R                      2                           2
--R   (39)  [quotient= (y  - 2y + 1)z,remainder= (- y  + 2y - 1)z]
--R   Type: Record(quotient: Polynomial(Integer),remainder: Polynomial(Integer))
--E 39
--S 40 of 46

qr.remainder
--R
--R
--R             2
--R   (40)  (- y  + 2y - 1)z
--R                                                    Type: Polynomial(Integer)
--E 40
--S 41 of 46
p - ((x+1) * qr.quotient + qr.remainder)
--R
--R
--R   (41)  0
--R                                                    Type: Polynomial(Integer)
--E 41
--S 42 of 46
p/q
--R
--R
--R         (y - 1)z
--R   (42)  --------
--R           z + 5
--R                                          Type: Fraction(Polynomial(Integer))
--E 42
--S 43 of 46
(2/3) * x**2 - y + 4/5
--R
--R
--R               2  2   4
--R   (43)  - y + - x  + -
--R               3      5
--R                                          Type: Polynomial(Fraction(Integer))
--E 43
--S 44 of 46
% :: FRAC POLY INT
--R
--R
--R                    2
--R         - 15y + 10x  + 12
--R   (44)  -----------------
--R                 15
--R                                          Type: Fraction(Polynomial(Integer))
--E 44
--S 45 of 46
% :: POLY FRAC INT
--R

--R
--R               2  2   4
--R   (45)  - y + - x  + -
--R               3      5
--R                                          Type: Polynomial(Fraction(Integer))
--E 45
--S 46 of 46
map(numeric,%)
--R
--R
--R                                            2
--R   (46)  - 1.0 y + 0.6666666666 6666666667 x  + 0.8
--R                                                      Type: Polynomial(Float)
--E 46
)spool
)lisp (bye)

———-

References^{[1] nothing}

12